# On the Viability of Open-Source Financial Rails: Economic Security of Permissionless Consensus

### v1

@@ Appendix D Other Factors Deterring Attacks
This section explores additional factors that deter attacks. Despite numerous attacks on cryptocurrencies1010todo: 10add list using the Nakamoto protocol, major cryptocurrencies such as Bitcoin and Ethereum have remained secure. The rental model and the bribery model discussed in Section 3 fail to explain why attacks on these protocols are not commonplace. This indicates that, in practice, other factors contribute to deterring attacks. We discuss such factors and identify elements that can be incorporated into protocol design to enhance security.
@@
In practice, the community can identify the attack fork and the honest fork. As described in Section 3, in the view of an honest node, an attack appears to be a reconciliation after a network partition. That is, an honest node will receive the blocks in the attack fork significantly later than blocks of comparable height in the honest fork. [^69] The attacker must keep hidden the blocks in the attack chain $({\sf B}^{\prime}_{n},\dots,{\sf B}^{\prime}_{n+L})$ until the merchant finalizes the block ${\sf B}_{n}$ (otherwise, the merchant will not finalize ${\sf B}_{n}$). Thus, an honest miner receives block ${\sf B}^{\prime}_{n}$ only after receiving block ${\sf B}_{n+k}$, which it receives significantly later than block ${\sf B}_{n}$. Recall that honest nodes mine conflicting blocks ${\sf B}_{n},{\sf B}^{\prime}_{n}$ only if these blocks are mined simultaneously (up to network delays). Any honest node active during the attack can recognize the attack chain as either a deliberate attack or a highly improbable event of partition reconciliation.
@@
An attacker may be deterred by the threat of legal prosecution. In practice, to cash out from an attack, the attacker will need to withdraw fiat money from some exchange. However, exchanges are subject to know-your-customer (KYC) rules that require them to know the identities of their customers.747474 There are many examples of hackers returning illicitly obtained funds; for example: https://www.reuters.com/technology/defi-platform-poly-network-reports-hacking-loses-estimated-600-million-2021-08-11/, accessed August 2024.

### v2

@@ Appendix D Other Factors Deterring Attacks
This section explores additional factors that deter attacks. Despite numerous attacks on cryptocurrencies88todo: 8add list using the Nakamoto protocol, major cryptocurrencies such as Bitcoin and Ethereum have remained secure. The rental model and the bribery model discussed in Section 3 fail to explain why attacks on these protocols are not commonplace. This indicates that, in practice, other factors contribute to deterring attacks. We discuss such factors and identify elements that can be incorporated into protocol design to enhance security.
@@
In practice, the community can identify the attack fork and the honest fork under the attack described in 3.2. As described in Section 3, in the view of an honest node, an attack appears to be a reconciliation after a network partition. That is, an honest node will receive the blocks in the attack fork significantly later than blocks of comparable height in the honest fork. [^76] The attacker must keep hidden the blocks in the attack chain $({\sf B}^{\prime}_{n},\dots,{\sf B}^{\prime}_{n+L})$ until the merchant finalizes the block ${\sf B}_{n}$ (otherwise, the merchant will not finalize ${\sf B}_{n}$). Thus, an honest miner receives block ${\sf B}^{\prime}_{n}$ only after receiving block ${\sf B}_{n+k}$, which it receives significantly later than block ${\sf B}_{n}$. Recall that honest nodes mine conflicting blocks ${\sf B}_{n},{\sf B}^{\prime}_{n}$ only if these blocks are mined simultaneously (up to network delays). Any honest node active during the attack can recognize the attack chain as either a deliberate attack or a highly improbable event of partition reconciliation.
@@
An attacker may be deterred by the threat of legal prosecution. In practice, to cash out from an attack, the attacker will need to withdraw fiat money from some exchange. However, exchanges are subject to know-your-customer (KYC) rules that require them to know the identities of their customers.818181 There are many examples of hackers returning illicitly obtained funds; for example: https://www.reuters.com/technology/defi-platform-poly-network-reports-hacking-loses-estimated-600-million-2021-08-11/, accessed August 2024.